OpenStudy (anonymous):

How many different arrangements can be made with the letters from the word MATH?

OpenStudy (anonymous):

4 choices for the first letter 3 for the second 2 for the third 1 for that last by the counting priniciple it is \[4\times 3\times 2\times 1\] also known as \(4!\)

OpenStudy (anonymous):

Thank you